OpenStudy (dtan5457):

At a card party, 2 door prizes are awarded by drawing 2 names randomly from a box of 40 names, including Joe and Pat. What's the probability that joe is not chosen for either prize?

OpenStudy (dtan5457):

Is this 39/40 x 38/39, 95%?

OpenStudy (dtan5457):

ANd what is the probability Pat wins 1 prize?

OpenStudy (dtan5457):

Is the 2nd one 1/40+1/39?

OpenStudy (dtan5457):

@Nnesha

OpenStudy (dtan5457):

@Loser66

OpenStudy (anonymous):

So there are 40 cards, only 2 are picked. 2 of them are two people. So Joe is one of them, what is the fraction?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

joe is not chosen for either prize.

OpenStudy (anonymous):

@Sup_Peeps

OpenStudy (xapproachesinfinity):

for a moment i thought 1-p would do

OpenStudy (dtan5457):

for which part?

OpenStudy (xapproachesinfinity):

prob that joe is not chosen

OpenStudy (dtan5457):

but they choose twice

OpenStudy (dtan5457):

@sleepyjess

OpenStudy (kropot72):

@dtan5457 Your solution to the first part (p = 0.95) is correct. The second part can be solved by using combinations: \[\large P(Pat\ wins)=\frac{39C1}{40C2}\]

OpenStudy (dtan5457):

They both are extremely close in terms of the answer, why combinations, though?

OpenStudy (kropot72):

The question deals with sampling without replacement, when a sample of size n is taken from a population containing a items of one particular type, and b of others. The probability that the sample contains exactly x of the particular type is given by the relevant term in the hypergeometric distribution. \[\large p(x)=\frac{\left(\begin{matrix}a \\ x\end{matrix}\right)\left(\begin{matrix}b \\ n-x\end{matrix}\right)}{\left(\begin{matrix}a+b \\ n\end{matrix}\right)}\]

OpenStudy (kropot72):

Using the formula for the first part we get: \[\large p(0)=\frac{1C0 \times 39C2}{40C2}=0.95\]